OpenStudy (anonymous):

Find the reference angle for -600 deg. answer=30 deg. Is this correct?

OpenStudy (anonymous):

reference angle: smallest equivalent angle x is a reference angle (or equivalent) of y if for: (x + 360n) = y , n is an integer [ a whole number ] 30 + 360n = -600 360n = -630 n = -1.75 not a whole number (like 1.0000, 2.0000, 3, 4, -5..) also true: x + 360n = -600 equivalent angle = x = -600 - 360n. so for any whole number you plug in n. x = -600 - 360(0) = -600 x = -600 - 360(-1) = -240 x = -600 - 360(-2) = 120 x = -600 - 360(-3) = 480 for n = -2, you get the reference angle 120.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)